OpenStudy (anonymous):

2. The Nut Barn requires 15 pounds of mixed walnuts and pecans from their wholesaler. Walnuts cost $2.26 per pound and pecans cost $1.96 per pound. How many pounds of pecans are needed if they only want to pay $2.18 per pound for the mixture?

OpenStudy (anonymous):

Can someone help me find the answer to this? It's so confusing.

OpenStudy (anonymous):

The weight of the mixed nuts is 15 pounds.\[W+P=15\]The next equation deals with the money.\[2.26*W+1.96*P=2.18*15\]Solve the two simultaneous equations for W and P, the final weight of the walnuts and pecans respectively. Verify that \[\{W=11 \text{ lbs},P=4 \text{ lbs}\} \]

OpenStudy (anonymous):

How do you solve for w and x? that's the part I get confused at.

OpenStudy (anonymous):

"How do you solve for W and P ?" From the first equation, W=15-P. Replace W in the second equation with the expression 15-P . The result is an equation in one unknown, P, that you should be able to solve. Plug the result for P into the first equation and calculate the value for W. If you are not comfortable with the symbols W and P as unknowns, then replace then with the symbols x and y respectively if you wish.

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

You're welcome and thank you for the medal.